OpenStudy (anonymous):

\[\int\limits_0^{1/\sqrt{3}} \sqrt{x+ \sqrt{x^2+1}}dx\]

hartnn (hartnn):

i would start with a traditional approach, x = tan y anyone with smarter solution?

hartnn (hartnn):

i think \(u = \sqrt{x+ \sqrt{x^2+1}}\) will work better! :)

OpenStudy (anonymous):

could you give further hints into the problem? we only have three minutes to solve these.

ganeshie8 (ganeshie8):

Let \(x=\sinh y\), the integral becomes \[ \int\limits_{..}^{..}\sqrt{\sinh y+ \cosh y}~ \cosh y\, dy = \int\limits_{..}^{..}\sqrt{e^y}~ \cosh y\, dy \]

ganeshie8 (ganeshie8):

replace \(\cosh y\) by \(\dfrac{e^y+e^{-y}}{2}\) and rest is trivial

OpenStudy (michele_laino):

please try this substitution: \[\large \sqrt {{x^2} + 1} = x + t\] where \(t\) is the new variable

OpenStudy (anonymous):

is that euler's substitution or something like that?

hartnn (hartnn):

ganesh, excellent thought! would substituting the limits be a challenge in your approach?!

ganeshie8 (ganeshie8):

\[x=0\implies y=\arcsin h (0) = 0\] \[x=1/\sqrt{3}\implies y=\arcsin h (1/\sqrt{3}) = \ln(1/\sqrt{3} + \sqrt{1/3+1}) = \ln(\sqrt{3})\]

ganeshie8 (ganeshie8):

\(x=\sinh y =\dfrac{e^y-e^{-y}}{2}\) \(e^{2y}-2xe^y-1=0 \) \(y = \arcsin h(x)= \ln(x+\sqrt{x^2+1})\)

OpenStudy (anonymous):

Great work!

ganeshie8 (ganeshie8):

Ofcourse this hyperbolic substitution saves time only if you're familiar with that identity

hartnn (hartnn):

if we know \(\sinh y = \ln (x+\sqrt{x^2+1 })\) then we could easily think of modifying our original integral function to \(\large \sqrt{e^{\arcsin h \: x}} \) in that case, the substitution y = arcsinh x or x = sinh y would become very obvious...